A 1500kg car is moving at 42.0 m/s. Determine the stopping distance if the brakes supply a force of 2000 newtons.
Answer #1

apply F = ma

a = 2000/1500 = 1.334 m/s^2

now use V^2 = 2aS

S = 42*42/(2*1.334)

S = 661.169 m
Answer #2

deceleration = 2000/1500 = 4/3

42^2 = 2as

So, distance = 661.5m
Answer #3

acceleration a = 2000/1500 = 4/3 m/s^2


v^2-u^2 = 2as


0 -42^2 = 2* -4/3 * s


stopping distance s = 2646 m
Answer #4

KE = 1/2mv^2 = 1323000 J

KE as you can see from the equation is related to the mass and the velocity. We're solving for the distance to get to a velocity of 0. This means our KE will also be 0. So, the total KE calculated will equal the total W necessary to bring the car to a stop.

Now we use the equation W = Fd

rearranged to find our required variable we have:

d = W/F = 1323000/2000

= 661 m.

So, you're probably going to kill that family in front of you that you're braking for. Physics is rough.